The area of ​​the winding loop is a square with a side length of 1.2 m and 10 turns and a current of 1.1 A. How much torque is caused by the Earth's magnetic field when the axis of symmetry of the winding (plane normal) is directed a) to the west? b) to the north?


The magnetic flux density of the Earth's magnetic field is 50 μT, the inclination is 73º and the declination is 0º.


1
Expert's answer
2021-09-06T06:54:24-0400

Gives

l=1.2m

n=10

i=1.1A

"\tau=M\times B"

"M=niA"

A="1.2\times1.2=1.44m^2"



"M=10\times1.1\times1.44=15.84Am"

"B=50\mu T"

"\tau=MBsin\theta"

Put value


"\tau=15.84\times 50 \times 10^{-6} sin73°=7.54\times 10^{-4}Am^2"